5. A point source emits a sound with an average power output of 80.0w. a) What is the intensity of the sound 3.00m away from the source? b) How far away from the source would you have to be to experience an intensity of 1.00x108 W/m2?
Answer #1

5A.

Intensity = P/A

I = P/(4*pi*r^2)

I = 80/(4*pi*3^2)

I = 0.707 W/m^2

5B.

I = P/(4*pi*r^2)

r = sqrt (P/(4*pi*I))

r = sqrt (80/(4*pi*10^-8))

r = 25231.32 m = 25.23 km